# Hierarchical Prompting Taxonomy: A Universal Evaluation Framework for Large Language Models

### v2

@@ Abstract
Assessing the effectiveness of large language models (LLMs) in addressing diverse tasks is essential for comprehending their strengths and weaknesses. Conventional evaluation techniques typically apply a single prompting strategy uniformly across datasets, not considering the varying degrees of task complexity. We introduce the Hierarchical Prompting Taxonomy (HPT), a taxonomy that employs a Hierarchical Prompt Framework (HPF) composed of five unique prompting strategies, arranged from the simplest to the most complex, to assess LLMs more precisely and to offer a clearer perspective. This taxonomy assigns a score, called the Hierarchical Prompting Score (HP-Score), to datasets as well as LLMs based on the rules of the taxonomy, providing a nuanced understanding of their ability to solve diverse tasks and offering a universal measure of task complexity. Additionally, we introduce the Adaptive Hierarchical Prompt framework, which automates the selection of appropriate prompting strategies for each task. This study compares manual and adaptive hierarchical prompt frameworks using four instruction-tuned LLMs, namely Llama 3 8B, Phi 3 3.8B, Mistral 7B, and Gemma 7B, across four datasets: BoolQ, CommonSenseQA (CSQA), IWSLT-2017 en-fr (IWSLT), and SamSum. Experiments demonstrate the effectiveness of HPT, providing a reliable way to compare different tasks and LLM capabilities. This paper leads to the development of a universal evaluation metric that can be used to evaluate both the complexity of the datasets and the capabilities of LLMs. The implementation of both manual HPF and adaptive HPF is publicly available [^1] GitHub repository consisting of the code implementation of this work: https://github.com/devichand579/HPT.
Hierarchical Prompting Taxonomy: A Universal Evaluation Framework for Large Language Models
Devichand Budagam1, Sankalp KJ2, Ashutosh Kumar3, Vinija Jain4, 5thanks: Work does not relate to position at Amazon, Aman Chadha4, 5thanks: Work does not relate to position at Amazon 1Indian Institute of Technology Kharagpur, India, 2AI Institute, University of South Carolina, USA 3Rochester Institute of Technology, USA, 4Amazon GenAI, USA, 5Stanford University, USA
@@ 1 Introduction
Large language models (LLMs) have revolutionized natural language processing (NLP), enabling significant advancements across a wide range of applications. However, evaluating the capabilities of LLMs across various datasets remains a challenging task. Traditional prompting strategies, which apply the same prompt to all dataset samples regardless of task complexity, often result in sub-optimal evaluations. This limitation highlights the need for more sophisticated evaluation methods that can adapt to varying levels of task complexity.
Figure 1: Hierarchical Prompt Framework (HPF) consists of five distinct prompting strategies, each tailored to different levels of task complexity, ensuring the appropriate prompt is selected for the given task.
Prompting plays a vital role in unlocking the full potential of LLMs. By designing specific input prompts, the LLM’s responses can be guided, significantly influencing the quality and relevance of the output. Effective prompting strategies have enhanced LLM performance on tasks ranging from simple question-answering to complex reasoning and problem-solving. Recent research has explored various approaches to prompting and reasoning evaluation in LLMs. Chain-of-Thought (CoT) prompting Wei et al. (2022b) elicits step-by-step reasoning, improving performance on complex tasks. Specializing smaller models Fu et al. (2023) and using large models as reasoning teachers Ho et al. (2022) have demonstrated the potential for enhancing reasoning capabilities. Progressive-hint prompting Zheng et al. (2023) has been proposed to conceptualize answer generation and guide LLMs toward correct responses. Metacognitive prompting Wang and Zhao (2024) incorporates self-aware evaluations to enhance understanding abilities.
We introduce the Hierarchical Prompting Taxonomy (HPT), a set of rules that employs a Hierarchical Prompt Framework (HPF) shown in Figure 1, a structured evaluation method addressing the need for more sophisticated assessment of LLM task-solving capabilities. This framework employs five distinct prompting strategies, each tailored to different levels of task complexity. By assigning a score called the HP-Score based on the LLM’s performance across these levels, this work provides a comprehensive assessment of its ability to solve diverse tasks.
Furthermore, the taxonomy incorporates an Adaptive Hierarchical Prompt framework, enabling an LLM, called prompt-selector to dynamically select the appropriate prompting strategy based on a given task complexity. This adaptive approach automates the evaluation process, also provides a better understanding of the judgment skills of LLMs in determining task complexity. The main contributions of this paper are as follows:
@@
    We propose the Hierarchical Prompting Taxonomy (HPT), a set of rules designed to establish a universal measure of task complexity for both datasets and LLMs.
@@
    Based on the rules of the HPT, this work introduces a set of five unique prompting strategies, arranged from simplest to most complex, to provide a more precise assessment of LLMs and offer clearer insights.
@@
    We introduce the Adaptive HPF that allows us to examine the judgement skills of LLMs by automating the evaluation process and dynamically selecting the right prompting approach based on task complexity.
The rest of the paper is organized as follows: Section 2 provides an overview of related work on prompting and evaluation in LLMs. Section 3 presents HPT and the frameworks associated with it in detail. Section 4 describes the experimental setup and results. Section 5 concludes the paper, Section 6 discusses the limitations and finally, Section 7 discusses the ethical impact of this work.
@@ 2 Related Work
The advent of LLMs has revolutionized NLP by demonstrating significant improvements in few-shot and zero-shot learning capabilities. Brown et al. (2020) introduced GPT-3, a 175 billion parameter autoregressive model, showcasing its ability to perform a wide range of tasks such as question-answering, reading comprehension, translation, and natural language inference without fine-tuning. This study highlighted the potential of very large models for in-context learning, while also identifying limitations in commonsense reasoning and specific comprehension tasks. Similarly, Liu et al. (2021) surveyed prompt-based learning, emphasizing the role of prompt engineering in leveraging pre-trained models for few-shot and zero-shot adaptation to new tasks with minimal labeled data.
Emergent abilities in LLMs, which appear suddenly at certain scale thresholds, have also been a topic of interest. Wei et al. (2022a) examined these abilities in few-shot prompting, discussing the underlying factors and implications for future scaling. Complementing this, Kojima et al. (2022) demonstrated that LLMs could exhibit multi-step reasoning capabilities in a zero-shot setting by simply modifying the prompt structure, thus highlighting their potential as general reasoning engines. Further exploring reasoning capabilities, Qiao et al. (2023) categorized various prompting methods and evaluated their effectiveness across different model scales and reasoning tasks, identifying key open questions for achieving robust and generalizable reasoning.
Several studies have focused on improving the interaction and decision-making abilities of LLMs through innovative prompting techniques. Lo et al. (2023) proposed ASH (Actor-Summarizer-Hierarchical) prompting, which enhances LLM performance on complex web navigation tasks by decomposing the process into summarization and action prediction. This hierarchical approach significantly outperformed previous methods on tasks with lengthy trajectories. Kong et al. (2024) introduced role-play prompting to improve zero-shot reasoning by constructing role-immersion interactions, which implicitly trigger chain-of-thought processes and enhance performance across diverse reasoning benchmarks.
The robustness and safety of LLM outputs have also been addressed in recent research. Wang et al. (2023) developed INDUST, a benchmark to evaluate LLM resistance to inductive instructions containing false premises, showing that self-critique prompting can help models handle such challenging instructions more effectively. Additionally, Chen et al. (2024) provided a comprehensive survey on prompt engineering techniques aimed at reducing hallucinations and enhancing the reliability of LLM outputs through advanced methods like chain-of-thought and tree-of-thought prompting.
To facilitate the evaluation and understanding of LLM capabilities, Zhu et al. (2024) introduced PromptBench, a unified library encompassing a variety of LLMs, datasets, evaluation protocols, and adversarial prompt attacks. This modular and extensible tool aims to support collaborative research and advance the comprehension of LLM strengths and weaknesses. Zhou et al. (2023b) proposed DYNAICL, a framework for efficient prompting that dynamically allocates in-context examples based on a meta-controller’s predictions, achieving better performance-efficiency trade-offs compared to uniform example allocation.
Chang et al. (2022) proposed SpeechPrompt, a framework that reformulates tasks as conditional generation problems for pre-trained generative spoken language models. This method demonstrated competitive performance in speech classification with fewer trainable parameters and potential in more challenging tasks like automatic speech recognition.
Lastly, Yao et al. (2023) introduced the Tree-of-Thoughts framework, enabling LLMs to deliberate over coherent text units and perform heuristic searches for complex reasoning tasks. This approach generalizes over chain-of-thought prompting and has shown significant performance improvements in tasks requiring planning and search, such as creative writing and problem-solving games.
@@
These works collectively highlight the advancements in leveraging LLMs through innovative prompting techniques, addressing their emergent abilities, reasoning capabilities, interaction strategies, robustness, and evaluation methodologies. Despite significant advancements, the current research on LLMs reveals several limitations, particularly in terms of prompt design, handling complex reasoning tasks, and evaluating model performance across diverse scenarios. The emergent abilities of LLMs, while promising, often lack predictability and control, and the robustness of these LLMs in the face of misleading prompts remains a concern.
To address these challenges, we propose the HPT and HPF for evaluating different LLMs. This approach aims to systematically categorize and assess the effectiveness of various prompting techniques based on task complexity, enabling a more nuanced understanding of LLM behaviors and performance. By introducing a structured framework for evaluation, this work seeks to enhance the reliability and applicability of LLMs, facilitating their deployment in increasingly complex and dynamic environments.
3 Methodology
@@
Figure 2: Hierarchical Prompting Taxonomy (HPT): A taxonomy designed to assess task complexity based on the criteria: Basic Recall and Reproduction, Understanding and Interpretation, Analysis and Reasoning, and Application of Knowledge and Reasoning.
3.1 Datasets and Models
Experiments were performed on BoolQ Clark et al. (2019), CommonSenseQA (CSQA) Talmor et al. (2019), IWSLT-2017 en-fr (IWSLT) Cettolo et al. (2017), and SamSum Gliwa et al. (2019). BoolQ consists of approximately 16,000 True/False questions with corresponding passages from Wikipedia, designed for binary classification in reading comprehension. CSQA contains around 12,000 multiple-choice questions sourced from ConceptNet, aimed at evaluating model’s commonsense reasoning abilities. IWSLT is a parallel corpus with tens of thousands of English-French sentence pairs from TED Talks, used for machine translation tasks. SamSum features around 16,000 human-generated chat logs with summaries, facilitating dialogue summarization model development. On the aforementioned datasets, the performance of the instruction-tuned LLMs Llama 3 8B AI@Meta (2024), Phi 3 3.8B Abdin et al. (2024), Mistral 7B Jiang et al. (2023), and Gemma 7B Team et al. (2024) were evaluated on their validation splits, with the exception of SamSum, which utilized the test split and nearly 2300 hours of computation was used for all experiments on the V100 GPU.
3.2 Hierarchical Prompting Taxonomy (HPT)
The proposed HPT shown in Figure 2 is founded on principles that are crucial for humans to address the majority of reasoning and interpretation tasks, taking into account the following criteria:
  • Basic Recall and Reproduction: Remember and reproduce factual information without interpretation or analysis.
  • Understanding and Interpretation: Comprehend and explain the meaning of information, summarizing or clarifying content.
  • Analysis and Reasoning: Break down complex information, understand relationships, and solve problems using logical reasoning.
  • Application of Knowledge and Execution: Apply knowledge in practical situations, execute multi-step processes, and solve complex tasks.
3.3 Hierarchical Prompt Framework (HPF)
The HPT employs an HPF that comprises five different prompting strategies. The complexity levels of these prompting strategies within the framework are determined by the taxonomy criteria, which judges actions such as the requirement for external knowledge, the capacity to solve subproblems, and the capacity to provide step-by-step reasoning that takes the task’s complexity into account. The five different prompting strategies that make up the HPF are :
  • Role Prompting Kong et al. (2024): Prompts that define a role for the LLM in solving the task.

### v4

@@ Abstract
Assessing the effectiveness of large language models (LLMs) in performing different tasks is crucial for understanding their strengths and weaknesses. This paper presents Hierarchical Prompting Taxonomy (HPT), grounded on human cognitive principles and designed to assess LLMs by examining the cognitive demands of various tasks. The HPT utilizes the Hierarchical Prompting Framework (HPF), which structures five unique prompting strategies in a hierarchical order based on their cognitive requirement on LLMs when compared to human mental capabilities. It assesses the complexity of tasks with the Hierarchical Prompting Index (HPI), which demonstrates the cognitive competencies of LLMs across diverse datasets and offers insights into the cognitive demands that datasets place on different LLMs. This approach enables a comprehensive evaluation of an LLM’s problem-solving abilities and the intricacy of a dataset, offering a standardized metric for task complexity. Extensive experiments with multiple datasets and LLMs show that HPF enhances LLM performance by 2%-63% compared to baseline performance, with GSM8k being the most cognitively complex task among reasoning and coding tasks with an average HPI of 3.20 confirming the effectiveness of HPT. To support future research and reproducibility in this domain, the implementations of HPT and HPF are available here.
Codehttps://github.com/devichand579/HPT
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing (NLP), enabling significant advancements in a wide range of applications. Conventional evaluation frameworks often apply a standard prompting approach to assess different LLMs, regardless of the complexity of the task, which may result in biased and suboptimal outcomes. Moreover, applying the same prompting approach across all samples within a dataset without considering each sample’s relative complexity adds to the unfair situation. To achieve a more balanced evaluation framework, it is essential to account for both the task-solving ability of LLMs and the varying cognitive complexities of the dataset samples. This limitation highlights the need for more sophisticated evaluation methods that can adapt to varying levels of task complexity.
Figure 1: Hierarchical Prompting Framework includes five distinct prompting strategies, each designed for different levels of task complexity to ensure the appropriate prompt is selected for the given task. A $\checkmark$ indicates task completion, while a $\times$ signifies task incompletion.
Within this study, complexity is defined as the cognitive demands associated with solving a task or the cognitive load introduced by a prompting strategy on LLMs. Henceforth, the term ”complexity” will be applied solely in this context. Task complexity, within the realm of human cognition, pertains to the cognitive requirements that a task imposes, which includes the diverse levels of mental effort necessary for processing, analyzing, and synthesizing information. According to Sweller (1988), tasks become more complex as they require greater cognitive resources, engaging working memory in more demanding processes such as reasoning and problem-solving. Similarly, Anderson et al. (2014) highlights that human cognitive abilities span a continuum from basic recall to higher-order thinking, with increasing difficulty correlating to tasks that demand analysis, synthesis, and evaluation. When applied to LLMs, the complexity of prompting strategies can be systematically evaluated by mapping them onto this human cognitive hierarchy. This alignment allows for an assessment of how LLMs perform tasks that reflect varying degrees of cognitive load, thereby providing a structured framework for understanding the cognitive demands associated with various tasks. By imposing this cognitive complexity framework on LLMs, this paper establishes a universal evaluation method, grounded in human cognitive principles, that enables more precise comparisons of model performance across tasks with varying levels of difficulty.
This paper introduces the HPT, a set of rules that maps the human cognitive principles for assessing the complexity of different prompting strategies. It employs the HPF shown in Figure 1, a prompt selection framework that selects the prompt with the optimal cognitive load on LLM required in solving the task. HPF enhances the interaction with LLMs, and improves performance across various tasks by ensuring prompts resonate with human cognitive principles. The main contributions of this paper are as follows:
@@
    Hierarchical Prompting Taxonomy (HPT): The paper proposes HPT, a set of rules that maps prompting strategies onto human cognitive principles, thereby enabling a universal measure of task complexity for LLMs.
@@
    Hierarchical Prompting Framework (HPF): The HPF is a framework designed to select the most effective prompt from five distinct prompting strategies, which optimizes the cognitive load on LLMs during task completion. This framework not only facilitates a more accurate evaluation of LLMs but also enhances their performance, delivering more transparent insights.
@@
    Hierarchical Prompting Index (HPI): HPI [^1] HPI can be quantitatively assessed to analyze the cognitive abilities of an LLM and the cognitive demands imposed by datasets on LLMs, as both factors are interchangeably related to the complexity of tasks. quantitatively evaluates the task complexity of LLMs across various datasets, offering insights into the cognitive demands that each task imposes on different LLMs.
HPF can be effectively compared to an ”Open Book” examination as shown in Figure 2, where questions represent tasks and textbooks serve as prompting strategies. In this analogy, the exam questions vary in complexity, from simple factual recall to intricate analytical problems, analogous to the tasks in HPT that are evaluated based on their cognitive demands. Similarly, textbooks provide structured guidance for solving these questions, much like the HPF organizes prompts in increasing levels of complexity to support LLMs. For example, a straightforward glossary lookup corresponds to a low-complexity task, while solving a multi-step analytical problem requiring synthesis of concepts represents a high-complexity task. The effort a student invests in answering a question mirrors the HPI, which measures the cognitive load placed on the LLM. Just as students perform better with structured resources like textbooks, LLMs improve with well-designed hierarchical prompting strategies, enabling them to tackle progressively complex tasks effectively.
Figure 2: Analogical framework comparing the HPF with ”Open Book” examination methodology. The diagram illustrates how HPF components (below) mirror traditional educational assessment elements (above), with parallel relationships between task complexity levels, resource utilization (prompts/textbooks), and performance metrics (HPI/student effort). This comparison demonstrates how LLM task complexity scales similarly to educational assessment complexity, from simple lookup tasks to complex synthesis problems
The remainder of the paper is structured as follows: Section 2 reviews the related work on prompting and evaluation in LLMs. Section 3 details the HPT and its associated frameworks. Section 4 outlines the experimental setup, results, and ablation studies. Section 5 concludes the paper. Section 6 discusses the ethical impact of the work.
@@ 2 Related Work
The advent of LLMs has revolutionized NLP by demonstrating significant improvements in few-shot and zero-shot learning capabilities. Brown et al. (2020) introduced GPT-3, a 175 billion parameter autoregressive model, showcasing its ability to perform a wide range of tasks such as question-answering, reading comprehension, translation, and natural language inference without fine-tuning. This study highlighted the potential of very large models for in-context learning while also identifying limitations in commonsense reasoning and specific comprehension tasks. Similarly, Liu et al. (2021) surveyed prompt-based learning, emphasizing the role of prompt engineering in leveraging pre-trained models for few-shot and zero-shot adaptation to new tasks with minimal labeled data.
2.1 Prompt Engineering
Prompting plays a vital role in unlocking the full potential of LLMs. By designing specific input prompts, the LLM’s responses can be guided, significantly influencing the quality and relevance of the output. Effective prompting strategies have enhanced LLM performance on tasks ranging from simple question-answering to complex reasoning and problem-solving. Recent research has explored various approaches to prompting and reasoning evaluation in LLMs. Chain-of-Thought (CoT) prompting (Wei et al. 2022b) elicits step-by-step reasoning, improving performance on complex tasks. Specializing smaller models (Fu et al. 2023) and using large models as reasoning teachers (Ho, Schmid, and Yun 2022) have demonstrated the potential for enhancing reasoning capabilities. Emergent abilities in LLMs, which appear suddenly at certain scale thresholds, have also been a topic of interest. Wei et al. (2022a) examined these abilities in few-shot prompting, discussing the underlying factors and implications for future scaling. Complementing this, Kojima et al. (2022) demonstrated that LLMs could exhibit multi-step reasoning capabilities in a zero-shot setting by simply modifying the prompt structure, thus highlighting their potential as general reasoning engines. Yao et al. (2023) introduced the Tree-of-Thoughts framework, enabling LLMs to deliberate over coherent text units and perform heuristic searches for complex reasoning tasks. This approach generalizes over chain-of-thought prompting and has shown significant performance improvements in tasks requiring planning and search, such as creative writing and problem-solving games. Kong et al. (2024) introduced role-play prompting to improve zero-shot reasoning by constructing role-immersion interactions, which implicitly trigger chain-of-thought processes and enhance performance across diverse reasoning benchmarks. Progressive-hint prompting (Zheng et al. 2023) has been proposed to conceptualize answer generation and guide LLMs toward correct responses. Metacognitive prompting (Wang and Zhao 2024) incorporates self-aware evaluations to enhance understanding abilities.
@@ 2.1 Prompt Engineering
These works collectively highlight the advancements in leveraging LLMs through innovative prompting techniques, addressing their emergent abilities, reasoning capabilities, interaction strategies, robustness, and evaluation methodologies. Despite significant advancements, the current LLM research reveals several limitations, particularly in terms of prompt design, handling complex reasoning tasks, and evaluating model performance across diverse scenarios. While promising, the emergent abilities of LLMs often lack predictability and control, and the robustness of these LLMs in the face of misleading prompts remains a concern.
2.2 Prompt Optimization and Selection
The challenge of optimizing prompts for LLMs has been addressed in several key studies, each contributing unique methodologies to enhance model performance and efficiency. Shen et al. (2023) introduce PFLAT, a metric utilizing flatness regularization to quantify prompt utility, which leads to improved results in classification tasks. Do et al. (2024) propose a structured three-step methodology that contains data clustering, prompt generation, and evaluation, effectively balancing generality and specificity in prompt selection. ProTeGi (Pryzant et al. 2023) offers a non-parametric approach inspired by gradient descent, leveraging natural language ”gradients” to iteratively refine prompts. Wang et al. (2024) present PromISe, which transforms prompt optimization into an explicit chain of thought, employing self-introspection and refinement techniques. Zhou et al. (2023b) proposed DYNAICL, a framework for efficient prompting that dynamically allocates in-context examples based on a meta-controller’s predictions, achieving better performance-efficiency trade-offs compared to uniform example allocation.
These studies aim to automate prompt design, moving away from traditional manual trial-and-error methods while emphasizing efficiency and scalability across various tasks and models. They report significant improvements in LLMs performance, with enhancements ranging from 5% to 31% across different benchmarks. This body of work underscores the increasing importance of prompt optimization and selection in unlocking the potential of LLMs and points toward future research avenues, such as exploring theoretical foundations, integrating multiple optimization techniques, and distinguishing between task-specific and general-purpose strategies.
2.3 Evaluation Benchmarks
To facilitate the evaluation and understanding of LLM capabilities, Zhu et al. (2024) introduced PromptBench, a unified library encompassing a variety of LLMs, datasets, evaluation protocols, and adversarial prompt attacks. This modular and extensible tool aims to support collaborative research and advance the comprehension of LLM strengths and weaknesses. Further exploring reasoning capabilities, Qiao et al. (2023) categorized various prompting methods and evaluated their effectiveness across different model scales and reasoning tasks, identifying key open questions for achieving robust and generalizable reasoning. (Wang et al. 2021) introduced a multi-task Benchmark for robustness Evaluation of LLMs extends the original GLUE (Wang et al. 2018) benchmark to assess model robustness against adversarial inputs. It incorporates perturbed versions of existing GLUE tasks, such as paraphrasing, negation, and noise, to test models’ abilities with challenging data. The study highlights that despite their success on clean datasets, state-of-the-art models often struggle with adversarial examples, underscoring the importance of robustness evaluations in model development.
3 Hierarchical Prompting Taxonomy
@@
Figure 3: Hierarchical Prompting Taxonomy: A taxonomy designed to assess the complexity of prompting strategies based on the criteria: Basic Recall and Reproduction, Understanding and Interpretation, Analysis and Reasoning, and Application of Knowledge and Reasoning.
3.1 Governing Rules
Figure 3 illustrates the HPT, a taxonomy that systematically reflects human cognitive functions as outlined in Bloom (1956). Each rule embodies complex cognitive processes based on established principles from learning and psychology.
  1. 1.
    Basic Recall and Reproduction: This reflects the fundamental cognitive process of remembering and reproducing factual information without analysis or interpretation, which involves mere recognition or retrieval of knowledge from memory (Anderson et al. 2014).
  2. 2.
    Understanding and Interpretation: This corresponds to the second cognitive rule of (Bloom 1956), where individuals must not only recall information but also explain it in their own words, summarize key points or clarify the meaning of content. This rule demands an intermediate cognitive load involving information processing rather than retrieving it.
  3. 3.
    Analysis and Reasoning: This aligns with the analysis stage of (Bloom 1956), which involves higher cognitive functions such as comparison, contrast, and deep understanding of the underlying principles. It is more complex than mere understanding because it requires examining structure and identifying patterns and connections.
  4. 4.
    Application of Knowledge and Execution: This mirrors the application and evaluation stages of (Bloom 1956), where individuals must not only understand and analyze but also use knowledge to perform multi-step tasks, solve complex problems, and execute decisions. It represents the most cognitively complex tasks, which require synthesis of information and practical decision-making, highlighting the critical leap from understanding theory to executing it in practice.
In HPT, the progression from basic recall to application of knowledge reflects increasing cognitive complexity, consistent with educational and cognitive frameworks, where more advanced cognitive processes build on foundational ones, demanding deeper engagement and mental effort.
3.2 Hierarchical Prompting Framework
The HPF consists of five prompting strategies, each assigned a complexity level. These levels are determined by the degree to which the strategies are shaped by the four principles of the HPT. The complexity levels of the prompting strategies are assigned based on human assessment of their relative cognitive loads over a set of 7 different tasks, guaranteeing that the cognitive abilities of LLMs are in harmony with those of humans. This approach enables the assessment of tasks in terms of their complexity and the cognitive load they impose on both humans and LLMs by utilizing HPI. Section 4.4 examines the hierarchical structure of the HPF in conjunction with the LLM-as-a-Judge framework, validating that the cognitive demands on LLMs can be aligned with those of humans.
The set of five prompting strategies were chosen from a diverse range of existing strategies to populate the framework, guided by a human judgment policy, prioritizing comprehensiveness in cognitive demands rather than the sheer number of strategies. See Appendix A for more details. Consequently, the HPF can be replicated or expanded with other relevant prompting strategies that exhibit similar cognitive demands, making the framework adaptable. The five prompting strategies, listed from least to most complex, are as follows:
  1. 1.
    Role Prompting (Kong et al. 2024): Prompts that specify the role of the LLM in task resolution represent the lowest level of complexity within the framework, as they exhibit minimal influence from the four principles of the HPT.
  2. 2.